x^2 - 9x - 38 = -9 by completing the square
Answer by checkley77(12844)

x^2 - 9x - 38 = -9
x^2-9x=-9+38
x^2-9x=29
x^2-9x+20.25=29+20.25
(x-4.5)^2=49.25
x-4.5=sqrt49.25
x-4.5=7.0178
x=7.0178+4.5
x=11.5178 ans.